# MADation: Face Morphing Attack Detection with Foundation Models

### v2

@@ Abstract
Despite the considerable performance improvements of face recognition algorithms in recent years, the same scientific advances responsible for this progress can also be used to create efficient ways to attack them, posing a threat to their secure deployment. Morphing attack detection (MAD) systems aim to detect a specific type of threat, morphing attacks, at an early stage, preventing them from being considered for verification in critical processes. Foundation models (FM) learn from extensive amounts of unlabeled data, achieving remarkable zero-shot generalization to unseen domains. Although this generalization capacity might be weak when dealing with domain-specific downstream tasks such as MAD, FMs can easily adapt to these settings while retaining the built-in knowledge acquired during pre-training. In this work, we recognize the potential of FMs to perform well in the MAD task when properly adapted to its specificities. To this end, we adapt FM CLIP architectures with LoRA weights while simultaneously training a classification header. The proposed framework, MADation surpasses our alternative FM and transformer-based frameworks and constitutes the first adaption of FMs to the MAD task. MADation presents competitive results with current MAD solutions in the literature and even surpasses them in several evaluation scenarios. To encourage reproducibility and facilitate further research in MAD, we publicly release the implementation of MADation at https://github.com/gurayozgur/MADation .
@@ 1 Introduction
The high focus of the research community on the study of deep learning techniques in recent years has led to the development of high-performing systems in several fields, including face recognition (FR) [5, 20]. However, the same scientific advances used to improve the recognition power of FR systems can also be used to create efficient ways to attack them [15, 23], posing a threat to their secure deployment. Morphing attacks (MA) constitute an example of such threats, as their generation process aims at incorporating features from more than one identity, resulting in a sample that can be verified by multiple people by the same FR system. When left undetected, these attacks can lead to several dangerous situations [13, 7, 56], such as allowing multiple people to pass border control with the same passport or letting a criminal travel under the identity of another person [36]. To address this problem, several morphing attack detection (MAD) systems have been proposed [27, 7, 21, 17, 38, 46]. MAD algorithms aim at distinguishing unaltered images (bona-fide samples) from MAs to identify malicious samples at an early stage and preventing them from being considered for verification in critical processes.
@@
Foundation models (FM) are large-scale networks that can be trained with self-supervised learning, which allows them to learn from unlabeled data. The fact that no labelling is required for FMs’ training samples highly simplifies the data acquisition task, allowing FMs to be trained in massive and diverse datasets. This training paradigm results in models that can efficiently generalize to a wide variety of assignments [4], making them particularly beneficial for fields that address several tasks, such as natural language processing (NLP) [6] and computer vision (CV) [34, 47, 40, 45]. Despite the recent attention given to FMs, their adaption to perform biometrics tasks is still very limited. While very recent works have used FMs to generate synthetic face images [41], perform iris segmentation [22] and FR [11], the utility of FMs for most biometrics fields is still highly underexplored. This literature gap should be carefully addressed, especially taking into account that biometrics tasks such as MAD may strongly benefit from FMs’ high generalization power, provided their efficient adaption to the domain specificities of the downstream task [11] as well as high generalizability when it comes to sub-domains [45], i.e. different morphing mechanisms in the MAD case.
@@
This work explores the potential of using FMs as the basis to the downstream MAD task. Given the domain-specific nature of the MAD task and knowing that FMs often underperform in specialized settings [53], we propose to adapt the pre-trained FM CLIP [45] to MAD with low-rank adaption (LoRA). This allows the FM to better align its feature space with the specificities of the downstream MAD task and still take advantage of the built-in knowledge acquired during pre-training, while simultaneously training a header to perform classification. This adaption paradigm corresponds to our proposed framework, MADation. We further evaluate whether MADation is properly taking advantage of the FMs’ properties by comparing it with alternative FM and transformer-based methods. We start by assessing the importance of adapting CLIP to the downstream MAD task by evaluating its zero-shot performance on this task (TI). Then, to verify whether LoRA adaption improves MAD performance, we use the FM as a frozen feature extractor. In this scenario, the FM’s feature space is not aligned with the specificities of MAD and only the classification layer is trained (FE). Finally, to ensure that the performance levels achieved by MADation are not only deriving from its architecture but are also dependent on the FM’s built-in knowledge acquired during pre-training, we assess the importance of the FM’s pre-trained weights by comparing MADation with models following the same architectures trained from scratch (ViT-FS). The developed experiments highlighted the efficiency of MADation when compared with the remaining analyzed methods, reducing the average EER achieved by ViT-FS and FE by 16.93 pp. and 8.10 pp., respectively, for CLIP ViT-L. Furthermore, MADation revealed competitive performance levels when compared with recent MAD solutions, highlighting FM’s potential in domain-specific tasks such as MAD.
@@ 2 Related Work
MAD: MAs are face images that result from the fusion of identity information belonging to two or more identities, allowing them to be simultaneously verifiable as belonging to all of them. An example of a MA can be found in Figure 1. Both human observers and FR systems are vulnerable to MAs [24, 49], leading to dangerous situations, such as multiple people being able to pass border control with the same passport [7]. To address the threat posed by MAs, several studies have proposed MAD systems [27, 7, 21, 17, 38, 46]. These systems can address the MAD task from two different perspectives, depending on the MAD operational scenario. Differential MAD solutions [12] are fed two samples simultaneously: a live capture of the individual claiming that the investigated image represents their identity and the investigated image itself. Although this approach is useful in scenarios such as border control, the fact that two images need to be compared to perform the detection limits its applicability in several scenarios [12], e.g. analyzing stand-alone documents. Hence, several studies have developed single-image MAD systems [27, 7, 21, 17, 38, 46, 31, 55], which can detect whether the investigated image is a morph based only on its characteristics. Ramachandra et al. [46] introduced a handcrafted-feature-based approach that extracts textural features across multiple scales and classifies them using collaborative representation. In [17], the authors deviated from the common binary classification of the whole investigated image by learning to classify each of its pixels (or pixel blocks) as bona-fide or MA. Fang et al. [21] proposed an unsupervised approach that used self-paced learning to assign smaller weights to suspicious samples, which generally correspond to MAs, allowing for training a robust autoencoder for anomaly detection even when the training data is polluted with MAs. Neto et al. [38] determined whether the analyzed sample contained two independent identities by separating its identity information into two orthogonal latent vectors. In [7], an autoencoder was trained on bona-fide samples to distill identity knowledge to a MAD system, following distinct distillation techniques for bona-fide and MAs. In [55], MAD was performed with ViT architectures, showing promising results. Ivanovska et al. [31] suggested using diffusion models to develop a self-supervised model that learns how to reconstruct bona-fide images from noisy inputs. As the model is trained on bona-fide samples alone, it is expected to lead to higher error rates when fed with MAs, making it possible to identify them through anomaly detection. Huber et al. [27] promoted the SYN-MAD 2022 competition on face MAD based on synthetic training data, presenting a comprehensive analysis of the results achieved by the seven submitted approaches. This work also focuses on single-image MAD due to its wider utility in real-world scenarios, whether they offer a live probe or not.
@@
Foundation Models: FMs contain a considerable amount of trainable parameters which allow them to be trained in large and diverse datasets. This intensive training leads to high adaptability, which is particularly useful in areas that deal with a wide range of tasks, such as CV [34, 45, 22, 40, 41, 11]. DINOv2 networks [40] is a series of self-supervised pre-trained visual models able to generate universal features that can be used to perform both image-level and pixel-level visual tasks. Kirillov et al. [34] proposed the Segment Anything Model (SAM) to perform image segmentation on a wide range of domains. SAM’s powerful generalization ability allows it to work on new image distributions and be competitive with prior supervised results when applying zero-shot learning. Radford et al. [45] proposed Contrastive Language-Image Pretraining (CLIP), a multimodal FM trained to interpret visual and textual inputs simultaneously, which allows it to effectively learn the correlation between images and their textual description.
@@
Although vision FMs’ massive training data makes them generalizable, they still show poor performance when applied to specific settings [53], leading to the need to adapt ViT networks to the desired downstream task to increase the performance achieved with zero-shot learning [9, 10, 25]. AdaptFormer [9] used two identical MLP branches instead of the MLP block in the transformer encoder. One of them replicates the original network to help maintain its properties, while the other allows for task-specific fine-tuning. ViT-Adapter [10] led to state-of-the-art (SOTA) COCO results for plain ViT networks, by combining the reconstruction of fine-grained multi-scale features with the introduction of image-related inductive biases into the FM. Hu et al. [25] inserted trainable rank decomposition matrices in the pre-trained FM, allowing for low-dimensional reparameterization. During the adaption process, the FM’s pre-trained weights are kept frozen and only these matrices’ weights are updated, allowing for an effective adaption with minimal computational cost. Using these matrices, also known as Low-Rank Adaptation (LoRA) layers, to adapt FMs has led to performance improvements in a broad range of tasks, such as capsule endoscopy diagnosis [54], plant phenotyping [8] and FR [11]. In this work, we select LoRA to adapt CLIP to the MAD task due to its promising results reported in the literature, namely in biometrics [11].
@@
Despite the growing attention given to FMs by the scientific community, their application in biometrics is still limited to a small set of recent works [22, 41, 11]. In [22], SAM [34] was fine-tuned to the iris segmentation task. In [41], FMs were used to generate facial images using identity-conditioned information. A recent work [11] used LoRA [25] to fine-tune DINOv2 [40] and CLIP [45] FMs to the FR task under several data availability settings. The experiments showed that FMs can be efficiently used in FR, especially in low data availability scenarios, where the proposed technique surpassed models trained from scratch.
@@
In this work, we contribute to the list of recent advancements in FMs’ application to biometrics by proposing a MAD solution based on FMs, MADation. In particular, we recognize the FMs’ generalization potential and their utility to tasks such as MAD when properly adapted to their domain-specific characteristics [11]. To this end, MADation incorporates LoRA layers in the analyzed FM while simultaneously training an extra classification layer.
@@ 3.1 Preliminary on CLIP
In this work, we use CLIP due to its capacity to generalize well to domain-specific downstream domains when properly adapted [11]. We assess CLIP’s abilities as a MAD model by evaluating its zero-shot learning performance (Section 3.3). To this end, CLIP is fed with image-text pairs where the text input corresponds to the possible image labels (‘face image morphing attack’ and ‘bona-fide presentation’ based on the ISO/IEC 20059 standard [29]). For the remaining approaches, only the image encoder is used and CLIP works as a feature extractor on top of which a classification layer is added, following the recent successful utilization of FMs in downstream tasks such as FR [11] and image segmentation [34]. These approaches include our proposed framework, MADation (Section 3.2), and two additional frameworks designed to assess the importance of the FM’s built-in knowledge and of adapting its feature space to the downstream domain (Section 3.3). When applicable (TI, FE, and MADation), CLIP was initialized with the pre-trained weights made publicly available in [45] [^1] https://github.com/OpenAI/CLIP.
@@ 3.2 MADation
Fine-Tuninig with LoRA: As mentioned in Section 2, the FMs’ drop in performance when facing domain-specific scenarios such as MAD can be addressed with ViT adapters [9, 10, 25]. In this work, we select LoRA [25] to adapt the selected FM due to its promising results reported in the literature, namely in biometrics [11]. LoRA employs a low-dimensional reparameterization strategy, demonstrating effectiveness comparable to training the full parameter space [1] while greatly reducing the number of parameters that need to be updated. In this method, the pre-trained weights of the FM, $W_{0}\in\mathbb{R}^{d\times k}$, are kept unchanged, and trainable rank-decomposition matrices are added within each layer of the transformer, enabling effective adaptation with minimal computational overhead. The low-rank decomposition introduced by LoRA updates $W_{0}$ as follows:
$\footnotesize W_{0}+\Delta W=W_{0}+\gamma_{r}BA,$ (1)
where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ are the trainable rank-decomposition matrices, with the rank $r<<min(d,k)$, and $\gamma_{r}$ is a scaling factor. Originally, $\gamma_{r}$ was defined as $\frac{\alpha}{r}$, but this formulation often causes gradient collapse as the rank $r$ increases, resulting in a lack of performance gains despite the use of additional trainable parameters for fine-tuning [32]. To address this issue, rank-stabilized LoRA (rsLoRA) [32] modifies the scaling factor to $\frac{\alpha}{\sqrt{r}}$, preventing gradient collapse and enabling better performance at higher ranks. For this reason, we opt to employ rsLoRA to fine-tune CLIP, setting $\gamma_{r}=\frac{\alpha}{\sqrt{r}}$. Since $\gamma_{r}$ and the original CLIP weights are kept frozen, only the matrices $A$ and $B$ are updated. Once the adaption process is complete, the final model weights are calculated as $W=W_{0}+\gamma_{r}BA$. As no extra parameters are added, this approach maintains the original model’s computational efficiency during inference.
@@
$\footnotesize\begin{gathered}Q_{i}=W_{i}^{q}x+\gamma_{r}B_{i}^{q}A_{i}^{q}x,\\ K_{i}=W_{i}^{k}x,\\ V_{i}=W_{i}^{v}x+\gamma_{r}B_{i}^{v}A_{i}^{v}x,\end{gathered}$ (2)
@@
$\footnotesize Attention(Q_{i},K_{i},V_{i})=Softmax\Big{(}\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{k}}}\Big{)}V_{i},$ (3)
@@
$\footnotesize Multihead(Q,V,K)=Concat(head_{1},...,head_{k})W^{0}.$ (4)
@@
Classification: Using LoRA to fine-tune CLIP, allows it to produce a final embedding space adapted to the downstream MAD task. In this scenario, the FM works as a feature extractor on top of which classification can be performed. Hence, an additional fully connected layer with two output neurons followed by a softmax layer is added on top of the FM, resulting in the complete detector architecture. This layer is trained along with fine-tuning the LoRA parameters, using the binary cross-entropy loss:
$\footnotesize L_{BCE}=-(y\,log(\tilde{y})+(1-y)\,log(1-\tilde{y})),$ (5)

### v3

@@ Abstract
Despite the considerable performance improvements of face recognition algorithms in recent years, the same scientific advances responsible for this progress can also be used to create efficient ways to attack them, posing a threat to their secure deployment. Morphing attack detection (MAD) systems aim to detect a specific type of threat, morphing attacks, at an early stage, preventing them from being considered for verification in critical processes. Foundation models (FM) learn from extensive amounts of unlabelled data, achieving remarkable zero-shot generalization to unseen domains. Although this generalization capacity might be weak when dealing with domain-specific downstream tasks such as MAD, FMs can easily adapt to these settings while retaining the built-in knowledge acquired during pre-training. In this work, we recognize the potential of FMs to perform well in the MAD task when properly adapted to its specificities. To this end, we adapt FM CLIP architectures with LoRA weights while simultaneously training a classification header. The proposed framework, MADation surpasses our alternative FM and transformer-based frameworks and constitutes the first adaption of FMs to the MAD task. MADation presents competitive results with current MAD solutions in the literature and even surpasses them in several evaluation scenarios. To encourage reproducibility and facilitate further research in MAD, we publicly release the implementation of MADation at https://github.com/gurayozgur/MADation
@@ 1 Introduction
The high focus of the research community on the study of deep learning techniques in recent years has led to the development of high-performing systems in several fields, including face recognition (FR) [5, 20]. However, the same scientific advances used to improve the recognition power of FR systems can also be used to create efficient ways to attack them [15, 23], posing a threat to their secure deployment. Morphing attacks (MA) constitute an example of such threats, as their generation process aims at incorporating features from more than one identity, resulting in a sample that can be verified by multiple people by the same FR system. When left undetected, these attacks can lead to several dangerous situations [13, 7, 56], such as allowing multiple people to pass border control with the same passport or letting a criminal travel under the identity of another person [36]. To address this problem, several morphing attack detection (MAD) systems have been proposed [27, 7, 21, 17, 38, 46]. MAD algorithms aim at distinguishing unaltered images (bona-fide samples) from MAs to identify malicious samples at an early stage and prevent them from being considered for verification in critical processes.
@@
Foundation models (FM) are large-scale networks that can be trained with self-supervised learning, which allows them to learn from unlabelled data. The fact that no labelling is required for FMs’ training samples highly simplifies the data acquisition task, allowing FMs to be trained in massive and diverse datasets. This training paradigm results in models that can efficiently generalize to a wide variety of assignments [4], making them particularly beneficial for fields that address several tasks, such as natural language processing (NLP) [6] and computer vision (CV) [34, 47, 40, 45]. Despite the recent attention given to FMs, their adaption to perform biometrics tasks is still very limited. While very recent works have used FMs to generate synthetic face images [41], perform iris segmentation [22] and FR [11], the utility of FMs for most biometrics fields is still highly under-explored. This literature gap should be carefully addressed, especially taking into account that biometrics tasks such as MAD may strongly benefit from FMs’ high generalization power, provided their efficient adaption to the domain specificities of the downstream task [11], and high generalizability when it comes to sub-domains [45], i.e. different morphing mechanisms in the MAD case.
@@
This work explores the potential of using FMs as the basis for the downstream MAD task. Given the domain-specific nature of the MAD task and knowing that FMs often underperform in specialized settings [53], we propose to adapt the pre-trained FM CLIP [45] to MAD with low-rank adaption (LoRA), while simultaneously training a header to perform classification. This allows the FM to better align its feature space with the specificities of the downstream MAD task and still take advantage of the built-in knowledge acquired during pre-training. This adaption paradigm corresponds to our proposed framework, MADation. We further evaluate whether MADation properly takes advantage of FMs’ properties by comparing it with alternative FM and transformer-based methods. We start by assessing the importance of adapting CLIP to the downstream MAD task by evaluating its zero-shot performance on this task (TI). Then, to verify whether LoRA adaption improves MAD performance, we use the FM as a frozen feature extractor. In this scenario, the FM’s feature space is not aligned with the specificities of MAD and only the classification layer is trained (FE). Finally, to ensure that MADation’s performance is not only deriving from its architecture but is also dependent on the FM’s built-in knowledge acquired during pre-training, we assess the importance of the FM’s pre-trained weights by comparing MADation with models following the same architectures trained from scratch (ViT-FS). The developed experiments highlight the efficiency of MADation compared with the remaining methods, reducing the average EER by 16.93 pp. and 8.10 pp. compared to ViT-FS and FE, respectively, for CLIP ViT-L. Furthermore, MADation revealed competitive performance levels compared with recent MAD solutions, highlighting FM’s potential in domain-specific tasks such as MAD.
@@ 2 Related Work
MAD: MAs are face images that result from the fusion of identity information belonging to two or more identities, allowing them to be simultaneously verifiable as belonging to all of them. An example of an MA can be found in Figure 1. Both human observers and FR systems are vulnerable to MAs [24, 49], leading to dangerous situations, such as multiple people being able to pass border control with the same passport [7]. To address the threat posed by MAs, several studies have proposed MAD systems [27, 7, 21, 17, 38, 46]. These systems can address the MAD task from two different perspectives, depending on the MAD operational scenario. Differential MAD solutions [12] are fed two samples simultaneously: a live capture of the individual claiming that the investigated image represents their identity and the investigated image itself. Although this approach is useful in scenarios such as border control, the fact that two images need to be compared to perform the detection limits its applicability in several scenarios [12], e.g. analysing stand-alone documents. Hence, several studies have developed single-image MAD systems [27, 7, 21, 17, 38, 46, 31, 55], which can detect whether the investigated image is a morph based only on its characteristics. Ramachandra et al. [46] introduced a handcrafted-feature-based approach that extracts textural features across multiple scales and classifies them using collaborative representation. [17] deviated from the common binary classification of the whole investigated image by learning to classify each of its pixels (or pixel blocks) as bona-fide or MA. Fang et al. [21] proposed an unsupervised approach that used self-paced learning to assign smaller weights to suspicious samples, which generally correspond to MAs, allowing for training a robust autoencoder for anomaly detection even when the training data is polluted with MAs. Neto et al. [38] determined whether the analyzed sample contained two independent identities by separating its identity information into two orthogonal latent vectors. [7] trained an autoencoder on bona-fide samples to distil identity knowledge to a MAD system, following distinct distillation techniques for bona-fide and MAs. [55] performed MAD with ViT architectures, showing promising results. [31] developed a self-supervised diffusion model that reconstructs bona-fide images from noisy inputs. As the model is trained on bona-fide samples alone, it leads to higher error rates when fed MAs, which can be identified through anomaly detection. [27] promoted the SYN-MAD 2022 competition on MAD based on synthetic training data, presenting a comprehensive analysis of the results of seven submitted approaches. This work also focuses on single-image MAD due to its wider utility in real-world scenarios, whether they offer a live probe or not.
@@
Foundation Models: FMs contain many trainable parameters, enabling FM to learn from large and diverse datasets. This intensive training leads to high adaptability, which is particularly useful in areas that deal with a wide range of tasks, such as CV [34, 45, 22, 40, 41, 11]. DINOv2 networks [40] is a series of self-supervised pre-trained visual models able to generate universal features that can be used to perform both image-level and pixel-level visual tasks. The Segment Anything Model (SAM) [34] can perform image segmentation on a wide range of domains. Contrastive Language-Image Pretraining (CLIP) [45] is a multimodal FM trained to interpret visual and textual inputs simultaneously, effectively learning the correlation between images and their textual description.
@@
Although vision FMs’ prove to be generalizable for several downstream tasks, they achieve less optimal performance when applied to some specific settings [53], which require adapting FMs to the desired downstream task [9, 10, 25]. AdaptFormer [9] used two identical MLP branches instead of the MLP block in the transformer encoder. One of them replicates the original network to help maintain its properties, while the other allows for task-specific fine-tuning. ViT-Adapter [10] led to state-of-the-art (SOTA) COCO results for plain ViT networks, by combining the reconstruction of fine-grained multi-scale features with the introduction of image-related inductive biases into the FM. Hu et al. [25] inserted trainable rank decomposition matrices in the pre-trained FM, allowing for low-dimensional reparametrization. During the adaption process, the FM’s pre-trained weights are kept frozen and only these matrices’ weights are updated, allowing for an effective adaption with minimal computational cost. Using these matrices, also known as Low-Rank Adaptation (LoRA) layers, to adapt FMs has led to performance improvements in a broad range of tasks, such as capsule endoscopy diagnosis [54], plant phenotyping [8] and FR [11]. In this work, we select LoRA to adapt CLIP to the MAD task due to its promising results reported in the literature, namely in biometrics [11].
@@
Despite the growing attention given to FMs by the community, their application in biometrics is still limited to a few recent works [22, 41, 11]. In [22], SAM [34] was fine-tuned to iris segmentation. [41] used FMs to generate identity-specific facial images. A recent work [11] used LoRA [25] to fine-tune DINOv2 [40] and CLIP [45] FMs to the FR task under several data availability settings. The experiments showed that FMs can be efficiently used in FR, especially in low data availability scenarios, where the proposed technique surpassed models trained from scratch.
@@
In this work, we contribute to the list of recent advancements in FMs’ application to biometrics by proposing a MAD solution based on FMs, MADation. In particular, we recognize the FMs’ generalization potential and their utility to tasks such as MAD when properly adapted to their domain-specific characteristics [11]. To this end, MADation incorporates LoRA layers in the analysed FM while simultaneously training an extra classification layer.
@@ 3.1 Preliminary on CLIP
In this work, we use CLIP due to its capacity to generalize well to domain-specific downstream tasks when properly adapted [11]. We assess CLIP’s abilities as a MAD model by evaluating its zero-shot learning performance (Section 3.3). To this end, CLIP is fed with image-text pairs where the text input corresponds to the possible image labels (‘face image morphing attack’ and ‘bona-fide presentation’ based on the ISO/IEC 20059 standard [29]). For the remaining approaches, only the image encoder is used and CLIP works as a feature extractor on top of which a classification layer is added, following the recent successful utilization of FMs in downstream tasks such as FR [11] and image segmentation [34]. These approaches include our proposed framework, MADation (Section 3.2), and two additional frameworks designed to assess the importance of the FM’s built-in knowledge and of adapting its feature space to the downstream domain (Section 3.3). When applicable (TI, FE, and MADation), CLIP was initialized with the pre-trained weights made publicly available in [45] [^1] https://github.com/OpenAI/CLIP.
@@ 3.2 MADation
Fine-Tuning with LoRA: As mentioned in Section 2, the FMs’ drop in performance when facing domain-specific scenarios such as MAD can be addressed with ViT adapters [9, 10, 25]. In this work, we select LoRA [25] to adapt the selected FM due to its promising results reported in the literature, namely in biometrics [11]. LoRA employs a low-dimensional reparametrization strategy, demonstrating effectiveness comparable to training the full parameter space [1] while greatly reducing the number of parameters that need to be updated. In this method, the pre-trained weights of the FM, $W_{0}\in\mathbb{R}^{d\times k}$, are kept unchanged, and trainable rank-decomposition matrices are added within each layer of the transformer, enabling effective adaptation with minimal computational overhead. The low-rank decomposition introduced by LoRA updates $W_{0}$ as follows:
$W_{0}+\Delta W=W_{0}+\gamma_{r}BA,$ (1)
where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ are the trainable rank-decomposition matrices, with the rank $r<<min(d,k)$, and $\gamma_{r}$ is a scaling factor. Originally, $\gamma_{r}$ was defined as $\frac{\alpha}{r}$, but this formulation often causes gradient collapse as the rank $r$ increases, resulting in a lack of performance gains despite the use of additional trainable parameters for fine-tuning [32]. To address this issue, rank-stabilized LoRA (rsLoRA) [32] modifies the scaling factor to $\frac{\alpha}{\sqrt{r}}$, preventing gradient collapse and enabling better performance at higher ranks. For this reason, we opt to employ rsLoRA to fine-tune CLIP, setting $\gamma_{r}=\frac{\alpha}{\sqrt{r}}$. Since $\gamma_{r}$ is a constant and the original CLIP weights are kept frozen, only the matrices $A$ and $B$ are updated. Once the adaption process is complete, the final model weights are calculated as $W=W_{0}+\gamma_{r}BA$. As no extra parameters are added, the original model’s computational efficiency during inference is maintained.
@@
$\begin{gathered}Q_{i}=W_{i}^{q}x+\gamma_{r}B_{i}^{q}A_{i}^{q}x,\\ K_{i}=W_{i}^{k}x,\\ V_{i}=W_{i}^{v}x+\gamma_{r}B_{i}^{v}A_{i}^{v}x,\end{gathered}$ (2)
@@
$Attention(Q_{i},K_{i},V_{i})=Softmax\Big{(}\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{k}}}\Big{)}V_{i}.$ (3)
@@
$Multihead(Q,V,K)=Concat(head_{1},...,head_{k})W^{0}.$ (4)
@@
Classification: Using LoRA to fine-tune CLIP allows it to produce a final embedding space adapted to the downstream MAD task. In this scenario, the FM works as a feature extractor on top of which classification can be performed. Hence, an additional fully connected layer with two output neurons followed by a softmax layer is added on top of the FM, resulting in the complete detector architecture. This layer is trained along with fine-tuning the LoRA parameters, using the binary cross-entropy loss:
$L_{BCE}=-(y\,log(\tilde{y})+(1-y)\,log(1-\tilde{y})),$ (5)